find the solution of IVP y +(a +1)y = e(6+1)t, y(0) = 0,4(0) = 2 using Laplace transform.a=0
Answer #1


Y+(atly = e(a+bt, yO=0, x(0)=2 a=0 we get yll+yret, Y0=0, Y (O) = 2. Taking Laplace transform on the equation, [LLY (4)}